Circular No. 6416

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

V1333 AQUILAE S. A. Ilovaisky and C. Chevalier, Observatoire de Haute-Provence,

communicate: “Observations of V1333 Aql, the optical counterpart of the recurrent x-ray transient Aquila X-1, have been obtained this year at Haute-Provence with the 1.2-m telescope (+ TK1024 CCD camera) on Feb. 28; Mar. 19, 20, 28, and 29; Apr. 4 and 5; May 27, 28, 29, and 30; and June 1, 6, and 7. The object remained in quiescence (V = 19.2) until May 29 UT, when it began to brighten, signaling the start of a new activity cycle. Images obtained on June 7.0 give V = 17.6, which is typical of the average levels seen in previous cycles. The last observed activity cycle started around 1995 July 17 (IAUC 6192) and ended sometime between Sept. 17 and Oct. 14, when the source was found again in quiescence. Observations at other wavelengths (particularly in x-rays, where weak emission should be detectable) are strongly encouraged.”

COMET C/1996 B1 (SZCZEPANSKI)

Ephemeris from orbital elements on MPC 27080:

1996 TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
June 6 9^{h}55.^{m}52 −25^{o}38.^{′}6 1.992 2.185 86.^{o}8 27.^{o}6 14.3
16 10 07.54 −27 28.8 2.182 2.280 82.3 26.2 14.9
26 10 20.44 −29 17.2 2.370 2.378 78.1 24.7 15.4
July 6 10 34.03 −31 05.4 2.556 2.476 74.0 23.2 15.9
16 10 48.21 −32 54.3 2.739 2.576 70.0 21.8 16.4
26 11 02.88 −34 44.7 2.919 2.675 66.2 20.3 16.9
Aug. 5 11 17.97 −36 36.5 3.094 2.776 62.5 18.9 17.3
15 11 33.45 −38 29.8 3.265 2.876 58.9 17.6 17.7
25 11 49.30 −40 24.5 3.431 2.976 55.5 16.2 18.2
Sept. 4 12 05.48 −42 20.2 3.590 3.077 52.2 15.0 18.5
14 12 22.00 −44 16.6 3.743 3.177 49.1 13.9 18.9
24 12 38.85 −46 13.5 3.888 3.276 46.3 12.8 19.3
Oct. 4 12 56.04 −48 10.2 4.024 3.376 43.9 11.9 19.6
14 13 13.56 −50 06.4 4.152 3.475 41.9 11.1 19.9
24 13 31.42 −52 01.8 4.271 3.573 40.5 10.4 20.2
Nov. 3 13 49.60 −53 55.8 4.379 3.672 39.8 10.0 20.5
13 14 08.10 −55 48.3 4.476 3.769 39.8  9.7 20.8
1996 June 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green